GlueX TOF Meeting, March 21, 2017

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, March 21, 2017
11:00 am EDT
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from the previous meeting
  3. Run Review
    1. PMT gain history
  4. Calibration Status
  5. PMT base test in Hall
  6. PMT replacement plan
  7. Counter reconfiguration
  8. NIM Paper (Paul)
  9. Action Item Recap



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2017 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Thomas Britton, Brad Cannon, Mark Ito (chair), Simon Taylor, Beni Zihlmann

There is a recording of this meeting on the BlueJeans site. Use your JLab credentials.

Review of minutes from the previous meeting

We went over the minutes from February 21.

  • Thomas reported that the lucite shield simulation needs to be re-run with a version after Richard's fix to HDGeant's electromagnetic background generation. In simulation, the shield was doing more harm than good whereas we saw a benefit in the test.
  • Simon continues to work on improvements in TOF-charged-track matching.
  • We noticed that several of the questions about the on-board amplifiers for the bases that were asked last time remain unanswered.

PMT gain history during the Spring 2017 Run

Beni looked a histograms generated as part of the overall calibration effort. The energy spectrum of pulses show clear minimum ionizing peaks and those can be fit, separately for individual channels and runs to give a relative gain history for each PMT in the TOF. For the plots and more details of the analysis, see his wiki page.

All channels show a similar trend: (1) a slow gain reduction over the course of the "100 nA" running of less than 10%, (2) a step down in gain near the transition from "100 nA" to "150 nA" running, and (3) a flattening after that transition. We were not able to come up with a good explanation for features (2) and (3). the trends are visible in all PMTs; those close to the beamline as well as those far away, and in both layers.

Sasha thought that perhaps the cause was degradation of the photocathode of the PMTs, Beni thought that increased temperature of the tubes at high rate may be a contributor to halting the decline in the corresponding period.

Calibration Status

Beni is in the middle of the task, analyzing the ROOT trees as they become available from skims Sean Dobbs is doing on Spring-17 data. It takes about 2 or 3 days to analyze a single run.

PMT base test in Hall

Paul Eugenio drove one of the counters from the mini-TOF up to JLab two weeks ago and it was placed on tables in the Hall. The run ended before we were able to get the counter mounted in front of the TOF and take data with beam. Still Beni managed to look at some pulses from a modified base, on one end of the counter, with a prototype amplifier. He sees a slight degradation in rise time. There were some difficulties getting the amplifier to work; it has been re-soldered multiple times in the past several weeks. Finally Vladimir Popov had to supply a new board. He is currently working on a new board design specific to our bases.

The counter remains in the Hall. Tom Carstens has re-arranged things to keep the counter out of the way of foot traffic.

New PMT from-the-factory specs

Sasha has tabulated measured specs of the recent batch of PMTs we purchased as spares from Hamamatsu. See his table for the numbers (old DocDB credentials). He notes that these tubes have significantly higher gain that those from the original PMT order, some by over a factor of two.

Testing is about to begin at FSU.

We did a quick mental tabulation of spare PMT's

  • New order: 30
    • at FSU: 25
    • at JLab: 5
  • Spares in the Hall: 3
  • mini-TOF: 10
    • on counters mounted on fixture at FSU: 8
    • on counter in Hall: 2

Total: 40

PMT replacement plan and counter reconfiguration

We discuss how to approach PMT replacement and/or PMT base modification (to add on-board amplifiers). Both operations require removal of the PMT from its lightguide, therefore if we do one then we should do the other at the same time. There is probably time to do that this summer, but we would need to start right away and so far we have not tested the high-rate performance of the modified bases under realistic conditions. In particular, we do not know how well the read-out will handle the DC baseline shift.

Beni's study of the gain history (above) tells us that the PMT are not on the way to near-term failure at the rates we ran in Spring-17.

Given all of this, we agreed it might be best to keep the current PMTs and bases as they are for the Fall-17 and Spring-17 runs. After that we can make the changes with more information about the amplifier design. And in fact we may be ready to do a physical re-configuration of the counters as Paul has proposed in the past. A final decision will require further discussion.

NIM Paper

In his absence, we decided to impose a deadline for a first draft of the NIM paper on Paul. After furious discussion we decided on the May Collaboration meeting as a reasonable due date (May 15).

Action Item Recap

  1. Write a draft of the NIM paper (Paul)
  2. Prepare study of scalar rates as a function of time for next meeting (Brad)
  3. Re-run Lucite shield simulation (Thomas)
  4. Continue TOF-track matching studies/improvements (Simon)
  5. Call a meeting with Tim, Eugene, and Paul on the PMT replacement plan (Mark)